Displaying reports 981-1000 of 1079.Go to page Start 46 47 48 49 50 51 52 53 End
Reports until 18:06, Tuesday 22 August 2023
Environmental Seismic
Sander Vermeulen - posted 18:06, Tuesday 22 August 2023 - last comment - 14:10, Wednesday 23 August 2023(11288)
Seismic noise measurements - amplifier noise

[Ian, Sander]

We attempted to measure seismic noise in various locations using the Wilcoxon accelerometer. We used an AlphaLab LNA10 (see photo) to amplify the signal from the accelerometer, and acquired the data downstream of the amplifer with the Moku:Go. 

The problem we found is that the amplifier seems to introduce noise in the measurement that obfuscates the signal from the accelerometer. This can be seen by comparing spectra obtained with the accelerometer connected through the amplifier vs. having just the amplifier connected (i.e. having nothing connected to the amplifier input). The measured spectra in the two cases are almost indistinguishable at amplifier gains of 10 and 100.

The attached spectra are for measurements with gain=1000, where the red traces (input 1) are the measurements of the channel with the amplifier, and the blue trace (input 2) is a reference measurement of an unconnected Moku input. The faded red trace is with the accelerometer connected through the amplifier, the bright red trace is with just the amplifier. The accelerometer was oriented vertically (z-direction).

Further indication that the LNA10 introduces signifcant noise is found by comparing the spectra here to those in entry 11284, which have very different frequency dependences. Those spectra were taken without an amplifier (I think?).

Images attached to this report
Comments related to this report
Ian MacMillan - 14:10, Wednesday 23 August 2023 (11291)

The spectra in entry 11284, was taken with the LNA10. We were certainly seeing more noise than we were on Friday but what I found to be so strange was that there we could not get consistent measurements. For example in the swing space we found that when the Moku Go was being run off my laptops power there was a massive 60Hz line but when we plugged the Moku in the line was gone. And it was not consistent. if we unplugged the Wilcoxon 731A then plugged it back in we would get a different reading. If we power cycled the Moku we would get a different reading. Nothing was consistent. I would think that this might be a bad electrical connection but we checked them all and they seem to be fine. It might be worth it to get a 2-pin MIL-C-5015 connector, which is what the Wilcoxon uses just to be sure but I'm not sure why we are getting such inconsistent measurements.

GQuEST General
Torrey Cullen - posted 14:07, Tuesday 22 August 2023 - last comment - 14:14, Tuesday 22 August 2023(11284)
Calibration of Seismometer for Swing Space

[Ian, Sander, Torrey]

We are trying to take seismic data for buildings on campus. To do this we have borrowed a Wilcoxon 731A from the Cryo Lab. To calibrate this device we took it to the 40m and took measurements next to their seismometer doen the X-arm on Friday around 6 pm.

We first, for the Wilcoxon, we used the channel from elog 17717 that was not being used. In the image of the panel in elog 17717 we tried channel 5, with the corisponding channel name C1:X01-MADC0_EPICS_CH28 but it was a slow channel.

After struggling for a few hours we asked Koji if there was a fast channel we could use, and he lent us channel 4 on that board The channel was C1:ALS-X_SLOW_SERVO1_IN1, the X, Y, Z channels used for the 40m's seis were C1:PEM-SEIS_EX_Y_IN1 etc.. We then placed the borrowed wilcoxon 731A accelerometer (manual here) next to the one at the 40m. We used the one located at the X End of the 40m. Data was taken using diaggui of the power spectra of our seismometer and the x, y, and z directions of the 40m's. In addition the cross power spectrum and coherence between our seis and x,y,z of the 40m's. Rough plots of all this data can be found in the attached files as well as all of the raw data. Note we also took a 200s time series of the data just in case.

This should allow us to take seismometer data for around the swing space. Before we do we should double check what any additional gains the channel we plugged into is obtaining before readout.

Images attached to this report
Comments related to this report
Torrey Cullen - 13:46, Tuesday 22 August 2023 (11285)

Update 30 seconds after posting this: Apparently the x end seismometer at the 40m is not calibrated to anything. This data will need to be retaken.

Ian MacMillan - 14:14, Tuesday 22 August 2023 (11286)

Apparently the seismometer that we calibrated against was not itself calibrated. It was advised that we use the seismometers at the vertex. I looked for them when we were origionally doing this. I will have to ask someone where they are.

We may take data today and do the caibration later.

How To General
Ian MacMillan - posted 11:49, Wednesday 16 August 2023 (11281)
Note on Pytest

Often our code uses pytest and Lee's version wield.pytest. If given a file with these tests and they don't run, add the attached files to the folder. They tell pytest what to do and how it handle it correctly.

to make a test make a function with a name that states with "T_". Then run the command:

 pytest .

or replace the period with the file name and it will run all the tests in that file. Us the code in post about the GQuEST Spectrum models to see an example.

Non-image files attached to this report
Controls Optimal Controls
Ian MacMillan - posted 11:36, Wednesday 16 August 2023 - last comment - 11:36, Wednesday 16 August 2023(11207)
Correcting the Flat FOM gain to always be 1

[Ian, Lee]

NOTE: this should have been posted on June 21 2023 as this is the date that it was writen, but it was in my drafts for some reason. I have added latex since then

In the H2 system the gain for the flat FOM (F2) is 1/F1_gain, where F1_gain is the gain of the BNS FOM (F1). This works fine for the H2 norm because that is made up of the composite RMS from both FOMs so as one FOM's gain increases and the others decreases they cancel out. However, this is not the case in the H_infinity case. The H_infinity norm is the maximum value for the frequency response from the Zinf input to the Flat FOM output this value must be less than or equal to the gamma that we are using. Currently we are searching for a gamma around 1. The equation for the frequency response along the  Zinf input to the Flat FOM output is:

\[ \left | \frac{K(\omega)P(\omega)}{1-K(\omega)P(\omega)} F_{\text{flat}}(\omega)\right |\leq \gamma \]

where \( K(\omega)\) is the controller response, \(P(\omega)\) is the plant response and \(F_{\text{flat}}(\omega)\) is the Flat FOM's response. If the environmental noise is large like at low frequencies then \( K(\omega)P(\omega) \) must be large to control it. Thus at low frequencies the above equation becomes,

\[ \lim_{KP \to \infty }\left | \frac{K(\omega)P(\omega)}{1-K(\omega)P(\omega)} F_{\text{flat}}(\omega)\right | \approx\left |F_{\text{flat}}\right | \leq \gamma \]

Thus if we use the F2_gain=1/F1_gain then as F1_gain gets very small F2_gain gets very large. When F2_gain is very large, the gamma we are searching for (around 1) is much less than the actual H_infinty norm which is on the order of the F2_gain and the solver fails. This is what was causing the problems with the RMS plot. 

We switched the F2 gain to be all ones and acquired attached plot.

Note: when this PDF is printed the second color bar has the same colors as the first color bar.


Improvements for code and plots for paper:

  1. Fix the color bar problem (noted above) for printing.
  2. Make the color bar discreet
  3. decrease the number of F1_gain points so that the plot is not so crowded 
  4. Maybe add an H2 boundary line. See how it looks
  5. change the F1_gain color bar. maybe connect like F1_gain points with a subtle line.
  6. Add an example of the model maker (getSPOFF) to Buzz
  7. make the code transpose the model right before solving in the H_infinity case
Non-image files attached to this report
Comments related to this report
Torrey Cullen - 10:54, Friday 23 June 2023 (11210)

Infinite tweaks made.

Seeking control perfection.

Efficiency thrives.

GQuEST General
Lee McCuller - posted 18:11, Monday 14 August 2023 - last comment - 13:31, Wednesday 16 August 2023(11280)
Geontropic Spectrum models

Here is python code as a wield.pytest that generates the Geontropic noise model from arXiv:2209.07543 [gr-qc, physics:hep-ph, physics:hep-th].

You'll need to install wield.utilities and wield.pytest to run. You can just drop  T_GQuEST_concept.py and GQuEST_models.py  in a folder and call pytest. The output is

Total power displacement 2.3306928419630585e-18
Effective_BW: 57.28 MHz  by c/L 0.954696717877494
Effective_BWsq: 14.68 MHz  by c/L 0.24468351704291535
Frequency of Max: 11.29MHz
3db_bandwidth: 12.70MHz
Total power displacement2 2.3306928419630585e-18
3db points:  [[ 5.49516136]
 [18.19782777]]
Effective_BW: 57.28 MHz  by c/L 0.954696717877494
Effective_BWsq: 14.68 MHz  by c/L 0.2446835170429154

GQuEST_model.pdf Shows the model with and without the IR cutoff. It also includes the integrated RMS. Note that the power spectrum falls off only as 1/F, so has a logarithmic divergence. We presume this logarithmic divergence gets cut of a c/w, for c the speed of light and w the beam radius. That's at about 100GHz for GQuEST's beam. That divergence is physically important, but not practically measurable.

GQuEST_model.pdf shows the same plot, but with overlayed axes to allow easy comparisons with the units of the Geontropic noise model paper. The code shows how to set up the plots for such fancy axes.

there are also plots showing the model in the Holometer.

 

Both of the plots also show that there is a reasonably close analytic formula for the noise. In python it is

Omega = 2*np.pi * F
Omega_SPEC = c_m_s/(2**0.5*L)
Canalytic = 2 * Cmax * Omega**2 * Omega_SPEC / (4 * Omega_SPEC**4 + Omega**4 - Omega**2 * Omega_SPEC**2)**0.5 / (Omega**2 + Omega_SPEC**2)**0.5

Which can be updated to be

\(\Omega = 2 \pi f\)

\(\newcommand{\obar}{\overline{\Omega}}\)\(\obar = \frac{c}{L}\)


\(S_L(f) \approx \frac{2 \Omega^2 \obar}{\sqrt{\Omega^4 - \frac{1}{2}\Omega^2 \obar^2 + \obar^4} \sqrt{2\Omega^2 + \obar^2}} S_L(f_{\text{pk}})\)

 

and Cmax is the maximum of the power spectrum. The above are in power, not amplitude, units.

I'll edit or add a comment with the Latex form once I have it.

 

Non-image files attached to this report
Comments related to this report
Ian MacMillan - 11:51, Wednesday 16 August 2023 (11282)

See note on pytest post

Also use matplotlib version 3.5.2. code does not graph on some of the most recent versions

Ian MacMillan - 13:31, Wednesday 16 August 2023 (11283)

The problem with matplotlib is in the wield utilities package. In wield.utilities mplfigb  line 420 and 421

The parameter "b" is deprecated in ax.grid (See Deprecations). It needs to be changed to "visible"

Lasers General (GQuEST)
Lee McCuller - posted 12:09, Thursday 10 August 2023 (11279)
Support manual for Boostik HP Amplifier

I requested the operating manual for our old-style Boostik HP amplifier. When we update the SOP's to include the amplifier, this should probably go with them.

 

Non-image files attached to this report
Lab Infrastructure General
Torrey Cullen - posted 14:44, Wednesday 09 August 2023 (11278)
Standard Operating Procedures for B102

https://dcc.ligo.org/M2300159

Rough draft for our swing space's SOP. We moved the fire extinguisher in the hall into the room. Just need the actual laser hazard sign out front now I believe and we can go laser hazard (no amplifier) in the swing space. Although we probably want to do some sort of cleaning before we expose optics to the dirty room.

Controls Optimal Controls
Ian MacMillan - posted 14:22, Tuesday 08 August 2023 - last comment - 14:07, Thursday 27 June 2024(11277)
Moving the Plant to Environmental Shaping Filter for Riccati Solver

In the Buzz code we move the plant to shape the seismic noise without shifting the controls noise is:

\[ S_c = \left |\frac{EP}{1-KP}\right |^2  S_\mathrm{env} + \left |\frac{MKP}{1-KP}\right |^2  S_\mathrm{meas}  \]

where \(K\) is the controller, \(P\) is the plant, \(E\) is the environmental noise shaping filter, \(M\) is the measurement noise shaping filter, \(S_c\) is the controls noise PSD, and  \(S_\mathrm{meas} \) and \(S_\mathrm{env} \) are the unshaped PSDs of the respective noise. If we move the plant to the env noise shaping filter such that \(E'=PE\),  \(P=1\), and  \(K'\) is a new controller calculated for this system then the controlls noise becomes

\[ S_c' = \left |\frac{EP}{1-K'}\right |^2  S_\mathrm{env} + \left |\frac{MK'}{1-K'}\right |^2  S_\mathrm{meas}  \]

In order for these two ststments to be equlivilant

\[ K'=KP \]

I think this means there needs to be some sort of elimination of the \(P\) from the \(K'\) since what we are looking for is \(K\) and not \(K'\).

Comments related to this report
Ian MacMillan - 14:07, Thursday 27 June 2024 (11685)

Assuming that \(P' \neq1 \) . i.e. the plant has some shape left in it, then to find the controller to use in the system use:

\[ K=\frac{K'P'}{P} ,\]

where \(K\) is the controller to use in the system, \(P \) is the origional unmodified plant, \(P' \) is the modified plant, and  \(K' \) is the controller found by buzz.

Lasers General
Torrey Cullen - posted 16:46, Friday 04 August 2023 (11276)
More cryolab cavity locking

I moved the beam profiler to in reflection of the cavity to see if there was a discernable difference between the REFL beam when the cavity is locked and unlocked. There was not (see attached, should also more into this new beam profiler software and better ways to record data). I have not recovered alignment of the TRANS beam on the beam profiler. Aaron lent me a camera to use in transmission but we suspect the power incident on them is most likely too low for the poor QE camera. It seems like the TRANS beam should be much easier to find than this as the beam profile and camera diode are quite large. The cavity still appears to be locking on the mode we want.

To combat this I increased the power from the laser to 2.2mW coming from the source (allthough this puts you close to the maximum range on the TEC controller). This corresponds to ~1.3mW incident on the cavity. The 1811 PD saturates at far lower than this so I have a .5 ND filter in front of the PD.

Aaron thinks (from the elog I believe) that they were able to get up to 7mW out of the laser in the past. May be worth looking into this, but currently the range on the TEC controller doesn't allow this.

Images attached to this report
GQuEST General (GQuEST)
Hudson Alexander Loughlin - posted 19:37, Thursday 03 August 2023 (11275)
PDH Lock on the CMTF Table with a Red Pitaya
We used a red pitaya to lock a Mephisto laser to a reference cavity. Once the TEM00 mode is within the Mephisto's piezo's scanning range, a python script will automatically lock the laser to the reference cavity. The system held lock for over 30 minutes and was only unlocked when we disconnected the computer from the red pitaya to run other tests. We successfully locked and unlocked the laser from the cavity several times and the locking script worked every time. The lock acquisition is nearly instantaneous upon initiating the python script. Chris also successfully ran the script and locked the cavity, and doing so doesn't require much know-how about the specific setup. The script automatically calculates the PID set point and the initial voltage to feed the piezo when initiating a lock. The feedback loop is implemented as just an integrator, and could certainly be tuned up further. We have not yet implemented slow feedback control to the laser's temperature, so the laser will eventually loose lock if its frequency drifts relative to the cavity's by more than about 80 MHz. I uploaded the locking script to a github repo for the lab. It should be generally useful for PDH locking optical cavities with FPGAs. Attached files show the curve fit used to calculate the PID set point and initial piezo voltage, the locked TEM00 mode on a camera, and a schematic of the controls scheme.
Images attached to this report
Non-image files attached to this report
Lasers General
Torrey Cullen - posted 15:21, Thursday 03 August 2023 (11274)
Steady lock achieved on west cavity

[Lee, Sander, Torrey]

We are able to lock on the cavity mode from my previous post. This is a dump of the moku/loop gain settings for future reference. 

Images attached to this report
Non-image files attached to this report
Lasers General
Torrey Cullen - posted 15:13, Wednesday 02 August 2023 (11273)
Cryo Cavity Alignments

[Ian, Torrey]

We found the beam in transmission of the cavity when it flashes. Alignment of the beam into the cavity has so far been unsuccessful. This screenshot is an example of the beam we see. So far the alignment process has been:

1) Scan the cavity at large amplitude with the current to locate the REFL dip.

2) Lower the scan amplitude and adjust the temperature according such that the REFL dip stays in the scan.

3) Repeat until a scan is no longer needed and manual park the cavity on resonance using the temperature.

4) Attempt to adjust the input mirrors while parked on resonance.

There a few problems with this. First, the smallest step size we can take on the temperature is too large. Makes it difficult to park on resonance. Second the stability of the laser's frequency isn't the best for this process. When attempted to manually park on resonance it will often drift out of range in a few seconds. The next logical step would be to lock the cavity on this mode, but we have been unsuccessful in doing so. Either the loop shape isn't quite right or the mode resonant in the cavity is just too garbage to lock on. 

 

Few more examples of things we saw: Aliasing effects, some garbage but was fixed due to aliasing effects of our scan frequency and the scanning slit frequency. Unsure what this is, we've seen it a couple of times but I'm not sure it is significant. Would be good to find out what kind of lenses are in the cavity chamber.

Images attached to this report
GQuEST General (GQuEST)
Hudson Alexander Loughlin - posted 09:00, Wednesday 02 August 2023 - last comment - 09:41, Wednesday 02 August 2023(11271)
GQuEST CMTF Laser Alignment to Cavity
Edoardo and I are aligning the laser in the CMTF to an optical cavity. We can get a lot of optical power into TEM01 and TEM10 modes, but can't manage to get power into a TEM00 mode, or find any spike corresponding to this mode on an oscilloscope when sweeping the laser's frequency by tuning its fast PZT input. We're driving the PZT from 0V to +100V and the laser's frequency shift per volt was previously measured to be 1.66 MHz/V (the spec in the manual is 1-2V/MHz). Calculating the frequency difference between the TEM00 and TEM01 modes, we find that it should be 150 MHz, so in fact we should not expect to see a TEM00 spike on the oscilloscope if the TEM01 mode is roughly centered on the scope, since our scanning range is only 165 MHz. To calculate this, see eqs. (10) and (11) in https://opg.optica.org/ao/abstract.cfm?uri=ao-23-17-2944 Given this information, we plan to adjust the laser temperature so the TEM01 mode is slightly out of the scanning range and then look for a spike corresponding to the TEM00 mode. We will update with our success or lack thereof.
Comments related to this report
Hudson Alexander Loughlin - 09:41, Wednesday 02 August 2023 (11272)GQuEST
This worked and the cavity is now well-aligned to a TEM00 mode.
Equipment General (GQuEST)
Daniel Grass - posted 23:31, Tuesday 01 August 2023 (11270)
Pump Quotes

Key: Pump (Pumping speed, minimum pressure, cost, size)

 

Agilent

Rougher:

IDP-3 (3.6 m^3/h, 0.25 Torr, $3610, 5.5 x 7 x 14 in) ($3,286 on Pro Vac)

IDP-7 (9.1 m^3/h, 0.02 Torr, $6,074 10 x 10 x 17 in)

IDP-10 (10.2 m^3/h, 0.02 Torr, $6,883, 10 x 12 x 17 in)

IDP-15 (15.4 m^3/h, 0.01 Torr, $8,705 , 13 x 14 x 19 in)

TriScroll 400 (15 m^3/h, 0.013 Torr, $8064, 11 in x ? x 14 in)

 

Turbo:

TwisTorr 74 (70 L/s, 4*10^-10 Torr, $5832, 2 kg)

TwisTorr 84 (80 L/s, 4*10^-10 Torr, $6252, 2 kg)

TwisTorr 305 (300 L/s, 8*10^-11 Torr, $11301, 5.8 kg)

TwisTorr 404 (400 L/s, 8*10^-11 Torr, $12727, 22.6 kg)

 

Integrated (seems like a bad deal unless they include a lot more than the individual pumps):

TPS-mini: TwisTorr 74 + 0.6 m^3/h diaphragm)

TPS-conpact with TwisTor 74 (3 m^3/h, $13944)

TPS-conpact with TwisTor 305 (3 m^3/h, $17,636)

TPS-flexy IDP-3 + TwisTorr 74: ($11937)

TPS-flexy IDP-3 + TwisTorr 305: ($17486)

TPS-flexy IDP-7 + TwisTorr 74: ($14423)

TPS-flexy IDP-7 + TwisTorr 84: ($15441)

TPS-flexy IDP-7 + TwisTorr 305: ($19594)

TPS-flexy IDP-10 + TwisTorr 305: ($19490)

 

Pfieffier (asked for quotes)

Rougher:

HiScroll 6: (6.1 m^3/h, 0.01 Torr, ?, 19 kg) ($5,888 from high vac depot)

HiScroll 12: (12.1 m^3/h, 0.004 Torr, ? 24 kg) ($7,709 from high vac depot)

HiScroll 18: (18.1 m^3/h, 0.004 Torr, ? 23 kg) ($9,144 from high vac depot)

MVP 030-3 (1.8 m^3/h, 1.5 Torr, ? 4.3 kg) (diaphragm)

MVP 070-3 (4.3 m^3/h, 0.75 Torr, ? 16.4 kg) (diaphragm) (Website doesn’t list under compatible with Turbos)

 

Turbo:

HiPace 10 (final pressure 4*10^-5 Torr)

HiPace 30 (32 L/s N2, 4*10^-10 Torr, ?, 3.2 kg)

HiPace 60 P (64 L/s N2, 7.5*10^-7 Torr, ?, 2.2 kg) *note bad pressure*

HiPace 80 (67 L/s N2, 4*10^-10 Torr, ?, 3.8 kg) ($6,555 from high vac depot)

HiPace 80 Neo (67 L/s N2, 8*10^-8 Torr, ?, 1.7 kg)

HiPace 300 (260 L/s N2, 4*10^-10 Torr, ?, 19.2 kg) ($13,110 from high vac depot)

ATH 500 (350 L/s N2, 8*10^-9 Torr, ?, 37.5 kg)

 

Integrated:

HiCube Eco backing pumps MVP 30 or smaller

HiCube 80 Classic

HiCube 80 Pro

 

Edwards

Rougher:

mXDS (3 m^3/h, 0.075 Torr, $4945, 8 kg)

nXDS (6-20 m^3/h, 0.005-0.022 Torr, $7,800-$14,000, 25kg) ($6000-$10,300 from high vac depot)

4 sizes

XDS (40 m^3/h, 0.02-0.06, $16,500-$18000, 48) ($11,692 from high vac depot)

 

Turbo:

nEXT55 (55 L/s N2, 1*10^-7 Torr, $8,174, ?)

nEXT85 (47-85 L/s N2, 4*10^-10 - 4*10^-9 Torr, $8,582, 2.9-4.4 kg) ($6,763 from high vac depot)

nEXT 240 (240 L/s, 4.5*10^-8 Torr, $13,195, 5.7 kg) ($11,000 from high vac depot)

nEXT 300 (300 L/s, 4.5*10^-8 Torr, $13,858, 5.7 kg)

nEXT730 (680 L/s N2, 1-6*10^-10 Torr, $20,742, 19.6 kg)

 

Integrated (don’t come with scroll)

T-Station 85

T-Station 300

 

 

Leybold

Rougher:

Screw Pump:

Ecodry 25 plus (25 m^3/h, 0.01 Torr, ?, ?) (I reached out for pricing)

 

Scroll Pumps:

ScrollVac 3 plus (3.5 m^3/h, 0.08 Torr, $4744, ?)

ScrollVac 3S plus (3.5 m^3/h, 0.08 Torr, $5219, ?)

ScrollVac 7 plus (36.1 m^3/h, 0.02 Torr, $7586, 26 kg)

Also have larger pumps

 

 

Iwata

Rougher

Scroll Meister

ISP-50 (3.6 m^3/h, 0.11 Torr, ?, 12 kg) ($4,835 from high vac depot)

ISP-90 (6.5 m^3/h, 0.037Torr, ?, 13 kg)

 

TMP-B70 (70 L/s N2, 10^-8 Torr, ?, 5.2 kg)

I reached out for pricing

 

I also checkedtequipment.net but nothing was cheaper than high vac depot

Lasers General
Torrey Cullen - posted 10:57, Tuesday 01 August 2023 - last comment - 12:55, Tuesday 01 August 2023(11262)
Cryolab Setup Diagrams and Lock Settings in multi-instrument mode

Attached are

1. The laser lock box settings via screenshot in multi-instrument mode using the PLL as the modulation setting.

2. Diagram of multi-instrument mode for reference.

3. Ariel photo of the set up.

4. Inkscape drawing of the set up.

Also should note when scanning the cavity using the current it should be safe to use the full range (2V). The TEC controller should be set to ~10.5kOhms, but double check the range on the integrator is roughly centered.

Images attached to this report
Non-image files attached to this report
Comments related to this report
Torrey Cullen - 11:03, Tuesday 01 August 2023 (11264)

https://imgur.com/a/udiuJuG

The log isn't letting me upload the photo of the lab. Uploaded to imgur instead.

Ian MacMillan - 12:55, Tuesday 01 August 2023 (11268)

I was able to upload the file here

Displaying reports 981-1000 of 1079.Go to page Start 46 47 48 49 50 51 52 53 End